# Tac-Man: Tactile-Informed Prior-Free Manipulation of Articulated Objects

### v3

@@ I Introduction
  • Ambiguity: Objects with similar external appearances may have vastly different internal structures, complicating the unique determination of their kinematics through perception. This issue is extensively explored by Zhu et al. [39].
@@
The paper is organized to provide a thorough exploration of Tac-Man and its implications. Sec. II reviews relevant literature, positioning our work within the broader context of robotic manipulation of articulated objects. Sec. III details Tac-Man, our novel tactile-informed, prior-free methodology. Sec. IV describes the setup and results of our real-world experiments, while Sec. V presents our simulation studies, showcasing the scalability and generalizability of Tac-Man. In Sec. VI, we present a detailed discussion on the strengths, limitations, and potential applications of Tac-Man, as well as considerations for future research. Finally, Sec. VII concludes the paper, summarizing our key findings and contributions to this field.
@@ II-D Contact Sensing
The effectiveness of Tac-Man relies on timely compensation for contact deviations, which are modeled geometrically as the changes in the contact point positions, effectively informing both the direction and the magnitude of the necessary adjustments. Under this formulation, sensors with a primary focus on sensing contact geometry, such as GelSight [62], TacTip [63], and others, can be directly adopted within the Tac-Man framework. Additionally, sensors primarily designed to measure contact forces can also be utilized within our Tac-Man framework, provided that they are properly calibrated according to Hooke’s law [66] to obtain geometric information. A detailed discussion on sensor selection is available in Sec. VI.
@@ II-E Tactile Servoing
Tactile servoing, which involves adjusting robot poses in response to tactile signals, has been a significant area of research since its inception by Weiss et al. [67] and first implementation by Berger and Khosla [68]. It has evolved into two main approaches: model-based and learning-based. Model-based tactile servoing uses explicit models such as the inverse tactile model [69], the tactile Jacobian model [70, 71], the projector matrix [72, 73], or other tactile-based controllers [74, 75] to map tactile feedback onto robot actions. Learning-based tactile servoing leverages modern machine learning techniques, including imitation learning [76] and deep learning [77, 78, 53], to enable implicit mapping for complex robot manipulations. Tac-Man builds upon the fundamental principles of model-based tactile servoing, demonstrating its efficacy in the intricate manipulation of articulated objects.
@@ III The Tac-Man Method
This section outlines our problem formulation for prior-free manipulation of articulated objects. At the core of Tac-Man is the ability to regulate stable contact throughout the manipulation process. This involves guiding the object from its initial to its final state using just tactile signals. We begin by introducing essential notations and preliminaries in Sec. III-A, as these provide the foundational basis for our problem formulation and the subsequent derivations detailed in Sec. III-B. We then describe the contact representation in Sec. III-C. This representation is crucial in determining whether a contact is stable and this is discussed in Sec. III-D. Finally, Sec. III-E presents a computational method for calculating the robot’s pose adjustments. These adjustments are essential for maintaining stable contact, thereby enabling the successful manipulation of articulated objects.
@@ III-A Notation and Preliminaries
[…]
@@ III-B Problem Formulation
$\boldsymbol{T}_{i+1}&=\boldsymbol{T}_{i}\boldsymbol{T}_{i+1}^{i}\\ &=\boldsymbol{T}_{i}\underbracket{\boldsymbol{T}_{r}^{i}}_{\text{execution}}\underbracket{\operatorname*{arg\,min}_{\boldsymbol{T}^{r}_{i+1}\in SE(3)}\left\|f_{c}(\boldsymbol{T}_{i}\boldsymbol{T}_{r}^{i}\boldsymbol{T}^{r}_{i+1},J)-\mathcal{C}_{1}\right\|_{2}}_{\text{recovery}}\end{aligned}.$ (6)
@@
The execution term in Eq. 6 involves proceeding to move along the preliminary direction as described in Eq. 4. This navigation progresses under the assumption of no prior knowledge about the object handle’s trajectory, posing a risk of exceeding the tolerance limits necessary for maintaining stable contact. To balance this risk and maintain efficiency, it is crucial to carefully select the magnitude of $t$ in Eq. 4. The objective is to maximize $t$ while ensuring that the constraints for stable contact are met. This optimization can be mathematically represented as:
[…]
@@
The recovery term in Eq. 6 implies that at each step, the robot adjusts its pose to minimize any deviation from the stable contact state $\mathcal{C}_{1}$, which may arise due to the execution step.
@@ III-C Contact Representation
For the effective computation of Eq. 6 and optimization of Eq. 7, a suitable representation of $\mathcal{C}_{i}$ is crucial. To ensure compatibility with the homogeneous transformation matrix and to match the capabilities of current tactile sensors, we represent the contact $\mathcal{C}_{i}$ as a set of positions of all discretized points at the contact site, relative to the reference frame $\{i\}$. The position of each contact point within $\mathcal{C}_{i}$ is denoted by $\boldsymbol{p}^{i}$. To facilitate alignment with the homogeneous transformation matrix, the column vector representation of $\boldsymbol{p}^{i}$ is defined in an augmented format as follows:
@@
For simplicity, we assume a flat gripper surface, a common assumption in current robotic grippers and easily achievable with tactile sensor fabrication. Consequently, $p_{x}^{0}$ is consistent across all points and set to 00, attainable through coordinate transformation. The position of a contact point at $\mathcal{C}_{i}$ is then characterized by the condition:
@@ III-D Stable Contact Formulation
Similarly, we formulate $f_{s}(\mathcal{C}_{i},\mathcal{C}_{0})$ and $s$ to address concerns related to slipping in Eq. 2. Practically, to prevent sliding, the friction $f$ between the object and the gripper should not surpass the maximum static friction $f_{M}$. In our model, the friction at $\mathcal{C}_{i}$ is positively correlated with the shear deformation observed from $\mathcal{C}_{0}$:
@@
Additionally, in the context of Eq. 3, we intuitively compute it as the difference between the corresponding points of two contacts, considering our established contact representation. To quantify this difference, we apply the loss function outlined in Besl et al. [79]. However, to ensure fairness in cases where a contact has a larger number of contact points, we normalize this loss by the number of corresponding pairs in $\mathcal{K}_{1i}$:
@@ III-E Robot Pose Update
Having established the contact representation and computation methods for stable contact constraints, we should address the optimization in Eq. 7 to compute the value of $t$ for the determination of $\boldsymbol{T}_{r}^{i}$. However, in the absence of knowledge about the handle’s trajectory $J$, determining $t$ explicitly remains challenging. Therefore, we adopt an exploration mechanism by continuously executing with small increments in $t$ until any term in the constraints of optimization Eq. 7 reaches its upper bound, scaled by a safety margin $\alpha$ ($10<\alpha<$ 1). This process is concluded as
@@
To recover the contact to within the specified bounds after the execution stage, we employ the Kabsch algorithm [80] to compute the optimal transformation $\boldsymbol{T}^{r}_{i+1}$ for the recovery term in Eq. 6:
@@ IV Real-world Experiments
To assess the effectiveness of Tac-Man, our proposed approach to prior-free manipulation of articulated objects, we conducted a series of real-world experiments. These were designed with meticulous attention to detail, as outlined in Sec. IV-A. They aim to thoroughly evaluate key constraints and determine essential hyper-parameters for the task, as discussed in Sec. III. Our experiments showcase the strengths of Tac-Man, particularly its robustness and adaptability under various conditions. These include scenarios where priors for object articulation are ambiguous (Sec. IV-B), imperfect (Sec. IV-C), unknown (Sec. IV-D), or obsolescent (Sec. IV-E), as previously discussed in Sec. I.
@@ IV-A Experimental Setup
Tac-Man is designed for broad compatibility with robotic systems capable of facilitating 6D movement in the gripper. In experiments, we utilized a Kinova Gen3 7DoF robotic arm equipped with a Robotiq 2F-85 gripper (Fig. 2). To assess the contact constraints as outlined in Sec. III, we integrated two GelSight-type tactile sensors into the gripper.
@@
To further enhance our ability to validate these constraints, we adopt GelSight-type tactile sensors, which effectively provide geometry information at the contact surface. The back of the deformable pad is rigidly attached to a rigid acrylic sheet. Each sensor was specially fabricated with a unique $7\times 7$ grid of labeled markers (Fig. 3) on the front of the pad. Without any surrounding structural geometry, these markers share the same mechanical characteristics, allowing for calibration using consistent parameters. This grid aids in accurately identifying point correspondences between two contacts. Although this modification slightly reduces the resolution and depth reconstruction accuracy as compared to conventional GelSight sensors, its effectiveness in practical scenarios and in validating the contact constraints detailed in Sec. III is notable.
@@
The determination of values for the defined parameters is crucial. The following outlines the chosen values, with a summary provided in Tab. I:
  • The threshold for identifying contact points, $\epsilon$ in Eq. 10, is set to be adaptable. This is due to the material’s flexibility, as shown in Fig. 4, which results in deformation not only in the contact region but also in the surrounding area upon contact with the object. This deformation relates to $\Delta N$ and the handle’s shape, among other factors. For adaptability, $\epsilon$ is defined as the minimal $\Delta N$ that ensures at least 8 contact points, exceeding the number of markers in a line (7 in our sensors), thus ensuring these points are non-colinear.
  • The normal elastic deformation limit, $e_{n}$ in Eq. 13, is set at $2~{}\mathrm{mm}$. This value is chosen based on observations that within this range, the system can revert to its initial state after releasing contact. It is important to note that this value represents a conservatively reduced normal elastic deformation for enforcing stricter constraints.
  • The parameter $\delta_{0}$ in Eq. 17 is linked to handle material properties. To set a lower bound, an experiment with a slippery Lego piece, depicted in Fig. 5(a), identifies the $\delta^{M}$ values at different $\Delta N$. This is done by tracking the peak shear displacement for each marker, as shown in Fig. 5(b). The results from multiple trials, illustrated in Fig. 5(c), lead to setting $\delta_{0}$ at $1.5$ for safety.
  • The shear elastic deformation limit, $e_{s}$ in Eq. 13, is calculated as $e_{n}\delta_{0}$. Though smaller than the actual one, this value is deemed acceptable for stricter constraint enforcement.
  • The upper bound for another stable contact constraint, $d$ in Eq. 3, is experimentally determined to be $0.4~{}\mathrm{mm}$, ensuring it does not exceed $e_{n}$, $e_{s}$, and $\delta_{0}$ during manipulation.
  • The movement magnitude in the execution stage, the small increment in $t$ established in Eq. 7, is set as $\epsilon\delta_{0}$ to preserve stable contact during execution.
@@
Figure 5: $\boldsymbol{\delta_{0}}$ value acquisition. The experiment to obtain hyper-parameter $\delta_{0}$, as defined in Eq. 17: (a) The experiment begins with the arm pulling backward after the gripper has firmly grasped a slippery Lego piece. (b) Multiple trials are conducted at different values of $\Delta N$. Three representative trials are visualized to illustrate the process. The results follow the friction model well and $\delta^{M}$ can be easily identified. The results align well with the friction model, allowing for the straightforward identification of $\delta^{M}$. (c) The $\delta_{0}$ obtained from these multiple trials is visualized, and a value of $1.5$ is selected for safety considerations.
@@ IV-B Manipulation under Ambiguous Priors
Figure 6: Manipulation under ambiguous priors. Tac-Man showcases its proficiency in addressing the challenge caused by ambiguous priors, characterized by objects that, while visually indistinguishable, feature distinct articulation mechanisms. Our evaluation involves four articulated objects that look identical but differ in articulation: (a) a cabinet with a revolute joint on the left; (b) another cabinet with a revolute joint on the right; (c) a cabinet with a revolute joint at the bottom; and (d) a drawer with a prismatic joint. In scenarios (e)–(h), corresponding to objects (a)–(d), Tac-Man starts from the same position and in the same direction. The initial backward pull by the arm reveals variations in contact loss due to the differing articulations. Upon reaching a certain deviation threshold, the system triggers an adjustment in the arm’s pose. This adjustment, optimized by Eq. 23, aims to recover stable contact and re-align to the correct interaction direction. The arm then continues its backward pull. Through this iterative process of adjustment and execution, each object, starting from the same initial state, is maneuvered to a distinct final state reflective of its unique articulation. For video demonstrations, refer to Supp Video S1.
Ambiguous visual priors pose a significant challenge in robotic manipulation, especially when objects with visually indistinguishable appearances harbor varying articulation mechanisms. This dilemma is discussed in Sec. II-A and exemplified by Zhu et al. [39]. To illustrate, Fig. 6(a)–(d) features four objects that, despite their identical appearances, are equipped with different articulations: a cabinet with a revolute joint on the left (a), another with the joint on the right (b), a third with a joint at the bottom (c), and a drawer with a prismatic joint (d). This diversity underscores the complexity of inferring object kinematics solely from visual information.
@@
Tac-Man effectively demonstrates proficient manipulation of distinct objects in all four cases, as shown in Fig. 6(e)–(h), each corresponding to the configurations in Fig. 6(a)–(d). Starting from the same initial state with stable contact $\mathcal{C}_{1}$, Tac-Man employs an iterative two-stage cycle, beginning with an identical preliminary direction $\boldsymbol{R}^{1}$. During the execution stage, tactile feedback is utilized to discern deviations from the correct interaction direction by observing changes in the contact points. For example, the contact in (e) exhibits a shear counterclockwise deviation, indicating a need for an additional clockwise rotation in the robot’s movement, and conversely for the shear clockwise deviation observed in (f). In the scenario depicted in (g), a deviation in the normal direction $\Delta N$ signals the requirement for an added rotation for recovery. Conversely, in (h), the contact remains stable as the arm’s movement is congruent with the correct direction. When a predetermined deviation threshold is reached, the system enters the recovery stage, employing the optimization outlined in Eq. 23 to compute the optimal arm pose that compensates for the observed deviation, thereby realigning the gripper to the correct interaction direction. Following this adjustment, the process resumes along the preliminary direction $\boldsymbol{R}^{i}$. Through continual refinement to preserve stable contact, the arm successfully manipulates the objects into their intended states, as depicted in the last images in Fig. 6.

### v4

@@ I Introduction
  • Ambiguity: Objects with similar external appearances may have vastly different internal structures, complicating the unique determination of their kinematics through perception. This issue is extensively explored by Zhu et al[39].
@@
The paper is organized to provide a thorough exploration of Tac-Man and its implications. Sec.˜II reviews relevant literature, positioning our work within the broader context of robotic manipulation of articulated objects. Sec.˜III details Tac-Man, our novel tactile-informed, prior-free methodology. Sec.˜IV describes the setup and results of our real-world experiments, while Sec.˜V presents our simulation studies, showcasing the scalability and generalizability of Tac-Man. In Sec.˜VI, we present a detailed discussion on the strengths, limitations, and potential applications of Tac-Man, as well as considerations for future research. Finally, Sec.˜VII concludes the paper, summarizing our key findings and contributions to this field.
@@ II-D Contact Sensing
The effectiveness of Tac-Man relies on timely compensation for contact deviations, which are modeled geometrically as the changes in the contact point positions, effectively informing both the direction and the magnitude of the necessary adjustments. Under this formulation, sensors with a primary focus on sensing contact geometry, such as GelSight [62], TacTip [63], and others, can be directly adopted within the Tac-Man framework. Additionally, sensors primarily designed to measure contact forces can also be utilized within our Tac-Man framework, provided that they are properly calibrated according to Hooke’s law [66] to obtain geometric information. A detailed discussion on sensor selection is available in Sec.˜VI.
@@ II-E Tactile Servoing
Tactile servoing, which involves adjusting robot poses in response to tactile signals, has been a significant area of research since its inception by Weiss et al[67] and first implementation by Berger and Khosla [68]. It has evolved into two main approaches: model-based and learning-based. Model-based tactile servoing uses explicit models such as the inverse tactile model [69], the tactile Jacobian model [70, 71], the projector matrix [72, 73], or other tactile-based controllers [74, 75] to map tactile feedback onto robot actions. Learning-based tactile servoing leverages modern machine learning techniques, including imitation learning [76] and deep learning [77, 78, 53], to enable implicit mapping for complex robot manipulations. Tac-Man builds upon the fundamental principles of model-based tactile servoing, demonstrating its efficacy in the intricate manipulation of articulated objects.
@@ III The Tac-Man Method
This section outlines our problem formulation for prior-free manipulation of articulated objects. At the core of Tac-Man is the ability to regulate stable contact throughout the manipulation process. This involves guiding the object from its initial to its final state using just tactile signals. We begin by introducing essential notations and preliminaries in Sec.˜III-A, as these provide the foundational basis for our problem formulation and the subsequent derivations detailed in Sec.˜III-B. We then describe the contact representation in Sec.˜III-C. This representation is crucial in determining whether a contact is stable and this is discussed in Sec.˜III-D. Finally, Sec.˜III-E presents a computational method for calculating the robot’s pose adjustments. These adjustments are essential for maintaining stable contact, thereby enabling the successful manipulation of articulated objects.
@@ III-A Notation and Preliminaries
  • Calligraphic font letters, such as $\mathcal{C}$, are used to denote sets. The notation $\lvert\cdot\rvert$ represents the cardinality, or number of elements, in a set. Specially, $\mathcal{K}_{ij}$ refers to a set of corresponding element pairs from sets $\mathcal{C}_{i}$ and $\mathcal{C}_{j}$, formulated as:
@@ III-B Problem Formulation
$\begin{aligned} \boldsymbol{T}_{i+1}&=\boldsymbol{T}_{i}\boldsymbol{T}_{i+1}^{i}\\ &=\boldsymbol{T}_{i}\underbracket{\boldsymbol{T}_{r}^{i}}_{\text{execution}}\underbracket{\operatorname*{arg\,min}_{\boldsymbol{T}^{r}_{i+1}\in SE(3)}\left\|f_{c}(\boldsymbol{T}_{i}\boldsymbol{T}_{r}^{i}\boldsymbol{T}^{r}_{i+1},J)-\mathcal{C}_{1}\right\|_{2}}_{\text{recovery}}\end{aligned}.$ (6)
@@
The execution term in Eq.˜6 involves proceeding to move along the preliminary direction as described in Eq.˜4. This navigation progresses under the assumption of no prior knowledge about the object handle’s trajectory, posing a risk of exceeding the tolerance limits necessary for maintaining stable contact. To balance this risk and maintain efficiency, it is crucial to carefully select the magnitude of $t$ in Eq.˜4. The objective is to maximize $t$ while ensuring that the constraints for stable contact are met. This optimization can be mathematically represented as:
$\underset{t}{\text{maximize}}$ $t$ (7)
@@
The recovery term in Eq.˜6 implies that at each step, the robot adjusts its pose to minimize any deviation from the stable contact state $\mathcal{C}_{1}$, which may arise due to the execution step.
@@ III-C Contact Representation
For the effective computation of Eq.˜6 and optimization of Eq.˜7, a suitable representation of $\mathcal{C}_{i}$ is crucial. To ensure compatibility with the homogeneous transformation matrix and to match the capabilities of current tactile sensors, we represent the contact $\mathcal{C}_{i}$ as a set of positions of all discretized points at the contact site, relative to the reference frame $\{i\}$. The position of each contact point within $\mathcal{C}_{i}$ is denoted by $\boldsymbol{p}^{i}$. To facilitate alignment with the homogeneous transformation matrix, the column vector representation of $\boldsymbol{p}^{i}$ is defined in an augmented format as follows:
@@
For simplicity, we assume a flat gripper surface, a common assumption in current robotic grippers and easily achievable with tactile sensor fabrication. Consequently, $p_{x}^{0}$ is consistent across all points and set to 0, attainable through coordinate transformation. The position of a contact point at $\mathcal{C}_{i}$ is then characterized by the condition:
@@ III-D Stable Contact Formulation
Similarly, we formulate $f_{s}(\mathcal{C}_{i},\mathcal{C}_{0})$ and $s$ to address concerns related to slipping in Eq.˜2. Practically, to prevent sliding, the friction $f$ between the object and the gripper should not surpass the maximum static friction $f_{M}$. In our model, the friction at $\mathcal{C}_{i}$ is positively correlated with the shear deformation observed from $\mathcal{C}_{0}$:
@@
Additionally, in the context of Eq.˜3, we intuitively compute it as the difference between the corresponding points of two contacts, considering our established contact representation. To quantify this difference, we apply the loss function outlined in Besl et al[79]. However, to ensure fairness in cases where a contact has a larger number of contact points, we normalize this loss by the number of corresponding pairs in $\mathcal{K}_{1i}$:
@@ III-E Robot Pose Update
Having established the contact representation and computation methods for stable contact constraints, we should address the optimization in Eq.˜7 to compute the value of $t$ for the determination of $\boldsymbol{T}_{r}^{i}$. However, in the absence of knowledge about the handle’s trajectory $J$, determining $t$ explicitly remains challenging. Therefore, we adopt an exploration mechanism by continuously executing with small increments in $t$ until any term in the constraints of optimization Eq.˜7 reaches its upper bound, scaled by a safety margin $\alpha$ ($0<\alpha<1$). This process is concluded as
@@
To recover the contact to within the specified bounds after the execution stage, we employ the Kabsch algorithm [80] to compute the optimal transformation $\boldsymbol{T}^{r}_{i+1}$ for the recovery term in Eq.˜6:
@@ IV Real-world Experiments
To assess the effectiveness of Tac-Man, our proposed approach to prior-free manipulation of articulated objects, we conducted a series of real-world experiments. These were designed with meticulous attention to detail, as outlined in Sec.˜IV-A. They aim to thoroughly evaluate key constraints and determine essential hyper-parameters for the task, as discussed in Sec.˜III. Our experiments showcase the strengths of Tac-Man, particularly its robustness and adaptability under various conditions. These include scenarios where priors for object articulation are ambiguous (Sec.˜IV-B), imperfect (Sec.˜IV-C), unknown (Sec.˜IV-D), or obsolescent (Sec.˜IV-E), as previously discussed in Sec.˜I.
@@ IV-A Experimental Setup
Tac-Man is designed for broad compatibility with robotic systems capable of facilitating 6D movement in the gripper. In experiments, we utilized a Kinova Gen3 7DoF robotic arm equipped with a Robotiq 2F-85 gripper (Fig.˜2). To assess the contact constraints as outlined in Sec.˜III, we integrated two GelSight-type tactile sensors into the gripper.
@@
To further enhance our ability to validate these constraints, we adopt GelSight-type tactile sensors, which effectively provide geometry information at the contact surface. The back of the deformable pad is rigidly attached to a rigid acrylic sheet. Each sensor was specially fabricated with a unique $7\times 7$ grid of labeled markers (Fig.˜3) on the front of the pad. Without any surrounding structural geometry, these markers share the same mechanical characteristics, allowing for calibration using consistent parameters. This grid aids in accurately identifying point correspondences between two contacts. Although this modification slightly reduces the resolution and depth reconstruction accuracy as compared to conventional GelSight sensors, its effectiveness in practical scenarios and in validating the contact constraints detailed in Sec.˜III is notable.
@@
The determination of values for the defined parameters is crucial. The following outlines the chosen values, with a summary provided in Tab.˜I:
  • The threshold for identifying contact points, $\epsilon$ in Eq.˜10, is set to be adaptable. This is due to the material’s flexibility, as shown in Fig.˜4, which results in deformation not only in the contact region but also in the surrounding area upon contact with the object. This deformation relates to $\Delta N$ and the handle’s shape, among other factors. For adaptability, $\epsilon$ is defined as the minimal $\Delta N$ that ensures at least 8 contact points, exceeding the number of markers in a line (7 in our sensors), thus ensuring these points are non-colinear.
  • The normal elastic deformation limit, $e_{n}$ in Eq.˜13, is set at $2~\mathrm{mm}$. This value is chosen based on observations that within this range, the system can revert to its initial state after releasing contact. It is important to note that this value represents a conservatively reduced normal elastic deformation for enforcing stricter constraints.
  • The parameter $\delta_{0}$ in Eq.˜17 is linked to handle material properties. To set a lower bound, an experiment with a slippery Lego piece, depicted in Fig.˜5(a), identifies the $\delta^{M}$ values at different $\Delta N$. This is done by tracking the peak shear displacement for each marker, as shown in Fig.˜5(b). The results from multiple trials, illustrated in Fig.˜5(c), lead to setting $\delta_{0}$ at $1.5$ for safety.
  • The shear elastic deformation limit, $e_{s}$ in Eq.˜13, is calculated as $e_{n}\delta_{0}$. Though smaller than the actual one, this value is deemed acceptable for stricter constraint enforcement.
  • The upper bound for another stable contact constraint, $d$ in Eq.˜3, is experimentally determined to be $0.4~\mathrm{mm}$, ensuring it does not exceed $e_{n}$, $e_{s}$, and $\delta_{0}$ during manipulation.
  • The movement magnitude in the execution stage, the small increment in $t$ established in Eq.˜7, is set as $\epsilon\delta_{0}$ to preserve stable contact during execution.
@@
Figure 5: $\boldsymbol{\delta_{0}}$ value acquisition. The experiment to obtain hyper-parameter $\delta_{0}$, as defined in Eq.˜17: (a) The experiment begins with the arm pulling backward after the gripper has firmly grasped a slippery Lego piece. (b) Multiple trials are conducted at different values of $\Delta N$. Three representative trials are visualized to illustrate the process. The results follow the friction model well and $\delta^{M}$ can be easily identified. The results align well with the friction model, allowing for the straightforward identification of $\delta^{M}$. (c) The $\delta_{0}$ obtained from these multiple trials is visualized, and a value of $1.5$ is selected for safety considerations.
@@ IV-B Manipulation under Ambiguous Priors
Figure 6: Manipulation under ambiguous priors. Tac-Man showcases its proficiency in addressing the challenge caused by ambiguous priors, characterized by objects that, while visually indistinguishable, feature distinct articulation mechanisms. Our evaluation involves four articulated objects that look identical but differ in articulation: (a) a cabinet with a revolute joint on the left; (b) another cabinet with a revolute joint on the right; (c) a cabinet with a revolute joint at the bottom; and (d) a drawer with a prismatic joint. In scenarios (e)–(h), corresponding to objects (a)–(d), Tac-Man starts from the same position and in the same direction. The initial backward pull by the arm reveals variations in contact loss due to the differing articulations. Upon reaching a certain deviation threshold, the system triggers an adjustment in the arm’s pose. This adjustment, optimized by Eq.˜23, aims to recover stable contact and re-align to the correct interaction direction. The arm then continues its backward pull. Through this iterative process of adjustment and execution, each object, starting from the same initial state, is maneuvered to a distinct final state reflective of its unique articulation. For video demonstrations, refer to Supp Video S1.
Ambiguous visual priors pose a significant challenge in robotic manipulation, especially when objects with visually indistinguishable appearances harbor varying articulation mechanisms. This dilemma is discussed in Sec.˜II-A and exemplified by Zhu et al[39]. To illustrate, Fig.˜6(a)–(d) features four objects that, despite their identical appearances, are equipped with different articulations: a cabinet with a revolute joint on the left (a), another with the joint on the right (b), a third with a joint at the bottom (c), and a drawer with a prismatic joint (d). This diversity underscores the complexity of inferring object kinematics solely from visual information.
@@
Tac-Man effectively demonstrates proficient manipulation of distinct objects in all four cases, as shown in Fig.˜6(e)–(h), each corresponding to the configurations in Fig.˜6(a)–(d). Starting from the same initial state with stable contact $\mathcal{C}_{1}$, Tac-Man employs an iterative two-stage cycle, beginning with an identical preliminary direction $\boldsymbol{R}^{1}$. During the execution stage, tactile feedback is utilized to discern deviations from the correct interaction direction by observing changes in the contact points. For example, the contact in (e) exhibits a shear counterclockwise deviation, indicating a need for an additional clockwise rotation in the robot’s movement, and conversely for the shear clockwise deviation observed in (f). In the scenario depicted in (g), a deviation in the normal direction $\Delta N$ signals the requirement for an added rotation for recovery. Conversely, in (h), the contact remains stable as the arm’s movement is congruent with the correct direction. When a predetermined deviation threshold is reached, the system enters the recovery stage, employing the optimization outlined in Eq.˜23 to compute the optimal arm pose that compensates for the observed deviation, thereby realigning the gripper to the correct interaction direction. Following this adjustment, the process resumes along the preliminary direction $\boldsymbol{R}^{i}$. Through continual refinement to preserve stable contact, the arm successfully manipulates the objects into their intended states, as depicted in the last images in Fig.˜6.